# Using Active Learning to Improve Quasar Identification for the DESI Spectra Processing Pipeline

### v1

@@ Abstract
The Dark Energy Spectroscopic Instrument (DESI) survey uses an automatic spectral classification pipeline to classify spectra. QuasarNET is a convolutional neural network used as part of this pipeline originally trained using data from the Baryon Oscillation Spectroscopic Survey (BOSS). In this paper we implement an active learning algorithm to optimally select spectra to use for training a new version of the QuasarNET weights file using only DESI data, specifically to improve classification accuracy. This active learning algorithm includes a novel outlier rejection step using a Self-Organizing Map to ensure we label spectra representative of the larger quasar sample observed in DESI. We perform two iterations of the active learning pipeline, assembling a final dataset of 5600 labeled spectra, a small subset of the approx 1.3 million quasar targets in DESI’s Data Release 1. When splitting the spectra into training and validation subsets we meet or exceed the previously trained weights file in completeness and purity calculated on the validation dataset with less than one tenth of the amount of training data. The new weights also more consistently classify objects in the same way when used on unlabeled data compared to the old weights file. In the process of improving QuasarNET’s classification accuracy we discovered a systemic error in QuasarNET’s redshift estimation and used our findings to improve our understanding of QuasarNET’s redshifts.
@@ 1 Introduction
The Dark Energy Spectroscopic Instrument (DESI) is an ongoing spectroscopic survey that aims to observe over 40 million objects over the course of five years, with the goal of mapping the expansion history of the universe [1, 2]. DESI is mounted on the Mayall 4-meter telescope at Kitt Peak National Observatory and will map approximately a third of the night sky over the course of the survey [3, 4]. DESI takes an exposure of up to 5000 spectra at a time, a significant increase in data collection over previous spectroscopic surveys [5, 6, 7]. To date [8] approximately one million spectra have been released publicly as the DESI Early Data Release (EDR) [9, 10], with Data Release 1 (DR1) forthcoming.
@@
QuasarNET originally consisted of four convolution layers followed by a fully connected layer. This fully connected layer was then connected to a series of fully connected layers of 26 outputs each. The number of fully connected layers is determined by how many quasar emission lines QuasarNET is trained on. Intrinsically, QuasarNET is a line-finder network, and its primary output from each output layer is an estimated position for the emission line associated with that fully connected layer. The first 13 outputs of each line-finder unit is a confidence value that the given emission line is present within that portion or “box” subset of the wavelength range, a coarse-grained estimate of the line position. The second 13 outputs are fine-grained estimates of line positions within each of the 13 boxes and are floating point values between -0.1 and 1.1, where 0 corresponds with the lower edge of the box and 1 corresponds with the upper edge. Classification of spectra using these outputs is done by selecting a confidence threshold and a number threshold and assigning a quasar classification to any spectra where more lines than the number threshold have confidence values above the confidence threshold. This means that classification only uses the first 13 outputs of each line-finder unit. QuasarNET also provides a redshift estimate for each spectrum, producing an independent redshift estimate for each emission line detected in the spectrum. Redshift estimation is done by using the combined estimated line position (coarse + fine) to estimate a central wavelength of an emission line, which is then converted to a redshift using the rest-frame wavelength associated with that line-finder unit. [21] provides a more in-depth explanation of how QuasarNET works.
@@
After outlining our modifications to QuasarNET, we outline the active learning methodology and its application to improving QuasarNET’s classification accuracy and purity when coupled with the aforementioned network changes. Ideally we would like a comparable amount of DESI data to train a new weights file to improve QuasarNET’s handling of DESI data. The sheer quantity of spectra observed by DESI, however, makes visual inspection of DESI data on a similar scale unfeasible, so by construction any training dataset comprised of DESI data will have significantly fewer spectra than that comprised of eBOSS data. To combat this reduced amount of training data we apply the active learning methodology to select an optimal set of data to be labeled that will provide the largest improvement in the network when trained. We couple this with a novel outlier rejection step designed to further ensure that the spectra we select to label provide adequate coverage of the possible data space. By doing this we are able to significantly improve QuasarNET’s performance on DESI data even when using a much smaller training dataset of approximately 6k spectra, only 10% of the previous training sample.
Our quasar dataset significantly improves the completeness of our classifications without sacrificing purity, which we demonstrate by splitting our selected sample into training and hold out validation sets. We then demonstrate the new QuasarNET’s performance on unlabeled data by comparing the classification and redshift estimates of repeated exposures of the same objects. In this test we further demonstrate the power of our new dataset with improved repeat classification accuracy, with QuasarNET more consistently classifying objects the same way in multiple exposures.
@@ 3 Active Learning and Dataset Selection
During the first few months of the DESI survey a selection of deep tiles were visually inspected to quantify the performance of the automatic DESI classification pipeline [27, 28]. These efforts visually inspected approximately 16000 galaxy targets of various subtypes and approximately 8500 quasar targets. We initially attempted to train QuasarNET using a subset of these spectra as a training dataset, keeping all the quasar specific tiles and a small subset of galaxies from the galaxy tiles. The small size of this dataset limited the performance of QuasarNET.
To further improve performance we decided to label additional DESI spectra to supplement those already labeled in prior VI campaigns. Since the amount of human effort to label and classify spectra is non-trivial, we want to minimize the amount of new spectra necessary to improve QuasarNET’s performance while maximizing its associated impact. This is especially true even though we use only the first few months of DESI data as a potential training set. This data will be released as part of DESI’s DR1, and represents the first few months of DESI main survey observations, over 3.6 million spectra, processed in the “Guadalupe” data reduction. In order to select which of these spectra to label and include in training we use the active learning algorithm. For an in depth and comprehensive exploration of active learning, its implementation, and its mathematics, we recommend [29].
Briefly, active learning uses an algorithm’s classification uncertainty to quantify which unlabeled data is most confusing to the network. This confusing data is then passed to a human expert who labels them for inclusion in a new training run. By continuously training on confusing data we can maximize the performance of the algorithm with the smallest amount of data, as we are continuously filling the network’s (or other algorithm’s) gaps in knowledge. When used with a classification algorithm active learning will tend to pick data points that lie along decision boundaries, i.e. those that have the highest uncertainty and thus the largest possibility of classification confusion.
@@
The output of a single iteration of the entire active learning algorithm is represented in Figure 3. We start with a potential dataset of 3.6 million spectra and apply the SOM based outlier rejection step to eliminate 15.5 thousand spectra from the pool of potential training data. We then apply a cut, removing repeat exposures of the same object to ensure that we have unique objects to visually inspect rather than multiple exposures of the same objects. This reduces the pool of spectra to just below 3 million unique objects. In Figure 3 we plot a histogram of the entropy values, defined in Eq. 3.2, of the remaining spectra.
A red line indicates the cutoff we selected, above which there are 1000 spectra that we request to be visually inspected and labeled. We select 1000 spectra to balance the impact the new data has on the network against the speed at which we can visually inspect spectra. The plot shows two notable features: first, there is a sharp peak at zero, representing a large set of spectra for which there is high degree of confidence and agreement in the ensemble classification. Second, there is a sharp dropoff near 1 corresponding to the point at which a spectrum lies at the 50% confusion boundary between two classes, with no classifications in the third. Above this point in the histogram are spectra that show significant confusion between all three classes, rather than being either confused between only two classes or highly confident in one and weakly confident in the others.
@@ 4 Results
We ran the active learning pipeline twice, selecting 1000 spectra at each step to visually inspect and label. At the initial step we use 3761 DESI spectra from quasar dense tiles (“Deep-field VI” from Table 1 of [27]) plus an additional 839 DESI spectra from non-quasar dense tiles (a subset of all spectra in Table 1 of [28]) which we include to ensure that the dataset is not overly quasar dense. These 4600 DESI spectra were visually inspected and the results validated in [27, 28]. We then augment this dataset with an additional 15000 spectra from the previous eBOSS training dataset. We are motivated to supplement the DESI-only dataset for the first pass because the DESI only dataset does not fully cover the possible quasar space. For the second active learning step we use only the originally labeled DESI spectra plus the additional 1000 labeled spectra from the first active learning step to train the networks, excluding the supplemental eBOSS data, as we now consider it representative of the DESI quasar observation space, and believe that the active learning algorithm will select spectra to fill any deficiencies.
@@
For the visual inspection steps in the active learning pipeline we require that two separate experts inspect each spectra, assigning an estimated redshift and spectral classification as well as a “quality” rating (on a scale of 0-4, with 4 the most confident) that quantifies how confident the inspector is in their labels. Each spectra is then inspected by a “merger” who resolves discrepancies in the two results and assigns a final quality rating. QuasarNET only uses spectra whose final quality is above 2.5. A more detailed description of this procedure is available in [27] and [28].
@@
We present the results of this two-step procedure in two subsections, in the first we report results using the full truth dataset which we split into training and validation subsets. In the second we test the final DESI-specific weights file, trained on the entirety of the truth table rather than a subset. The weights testing used the internal data release “Guadalupe”, which will be included in DESI’s Data Release 1 (DR1), where we have no truth data but for which we do have a set of metrics that quantify this weights file’s improvement over the previous version. These metrics include consistency in classification and redshift uncertainty. In the second section we will also outline the new understanding of QuasarNET’s redshift estimation problems, which became apparent when using this newly trained weights file.
@@ 4.1 Results on Labeled Data
In this section we train a weights file using a subset of the full truth table, which includes the original set of visually inspected quasar tiles (3761 spectra), the additional subset of other galaxy dense visually inspected tiles (839 spectra), and the results of two active learning runs (2100 spectra). The total number of spectra in the truth table is 6700, although after training we discovered 2 spectra whose data processing was corrupted or otherwise nonphysical that were affecting training and removed them for a grand total of 6698 labeled spectra. We split this dataset into two subsets such that 70% of the data is used for training and the other 30% is used for testing and validation. In order to validate that the inclusion of the active learning spectra has improved the network’s performance notably we compare this training to two other training schemes. In one run we use the subset of the training dataset that excludes those spectra added after the second active learning step (keeping those spectra labeled in the first active learning run) and in the other run we use the same eBOSS dataset for training that was used to train the prior DESI weights file. All three runs use the same validation dataset. The results of these three runs are shown in Figure 4, with the eBOSS run in green, the results after the first visual inspection run in red and the results after both visual inspection runs in blue. For this plot, we use a confidence threshold of 0.95, thus declaring that QuasarNET identifies a quasar if it has at least one line with a confidence of more than 0.95.
@@
$I=1-P=1-\frac{N_{\text{QSO,TRUE \& QSO,PRED}}}{N_{\text{QSO,PRED}}}.$ (4.2)
As validated in the plot, the second visual inspection run gave a significant improvement in completeness at approximately the same (although marginally better) purity at each epoch, reaching a high maximum completeness faster than after the first visual inspection run. Also notable is the equivalent or higher completeness and purity per iteration at its maximum compared to the prior eBOSS training, validating that a DESI only dataset improves over an eBOSS one when used with DESI validation data.
@@
It is possible that a larger dataset could explain some or all of the improvement at each active learning step. In order to validate that our improvement is more significant than dataset size alone, we bootstrap each of the two earlier runs, the pre-active learning and the post first iteration active learning, to have the same amount of data as the final assembled dataset. We perform 10 unique bootstraps of each dataset, for 20 total unique datasets of the same size as the final dataset, and then train for 100 epochs with the same parameters as the previous test. The results of this test are shown in Figure 5. For clarity we plot only the completeness, although purity follows a similar trend. The solid lines represent the same completeness curves as plotted in Figure 4. In dashed blue and dashed magenta are the 10 bootstrap runs for the pre-active learning and post first iteration datasets, respectively. Although the increased amount of data does have some effect on the training it does not significantly alter the final best completeness values. In none of the bootstrap datasets does the training reach the level of completeness that our final dataset provides. This further validates that our active learning pipeline has specifically chosen spectra that give significant improvements to the training cycle.
@@
Figure 6: Plot of completeness versus purity for a variety of confidence thresholds ranging from zero to one. At each confidence threshold we compute quasar classifications and the corresponding completeness and purity on the validation dataset given those classifications. Colors match those of Fig. 4. A large dot on each curve (below the bottom of the plot for the original DESI visual inspection) denotes the values at the standard 0.95 confidence threshold. At a confidence threshold of 0.95, our new training post-active-learning outperforms the eBOSS weights file in completeness with a similar purity, although there are some threshold values at which the opposite is true.
@@
For our active learning pipeline we used a confidence threshold of 0.95, meaning that any spectrum with at least one line with a confidence more than 0.95 was considered a quasar. This motivated the choice used in the loss curves in Figure 4 and Figure 5, where the purity and completeness per epoch were calculated using this same confidence threshold. It is also informative to explore how the choice of confidence threshold changes the final results of the network. In Figure 6 we plot the purity and completeness against each other, with each point computed using a different confidence threshold between zero and one. This curve shows the trade off between improved completeness and lower purity at lower thresholds and the opposite at higher thresholds. Large dots indicate the values of both quantitites at our test threshold of 0.95. In the plot we see that at our threshold of 0.95 we have succeeded in improving the completeness of the final weights file, with almost identical purity. This effect does not hold true at all confidence thresholds. At lower thresholds the eBOSS weights file exceeds our newly trained weights file in completeness with similar or better purity. This is partially due to our use of this confidence threshold when determining classification for active learning, as we only pick spectra that are confused between classes when the constituent networks of the ensemble are very confident in their own classifications. It is likely that if we had used a lower threshold we would see our network outperforming the eBOSS one at that specific threshold, as we do with a threshold of 0.95.

### v2

@@ Abstract
The Dark Energy Spectroscopic Instrument (DESI) survey uses an automatic spectral classification pipeline to classify spectra. QuasarNET is a convolutional neural network used as part of this pipeline originally trained using data from the Baryon Oscillation Spectroscopic Survey (BOSS). In this paper we implement an active learning algorithm to optimally select spectra to use for training a new version of the QuasarNET weights file using only DESI data, with the goal of improving classification accuracy. This active learning algorithm includes a novel outlier rejection step using a Self-Organizing Map to ensure we label spectra representative of the larger quasar sample observed in DESI. We perform two iterations of the active learning pipeline, assembling a final dataset of 5600 labeled spectra, a small subset of the approximately 1.3 million quasar targets in DESI’s Data Release 1. When splitting the spectra into training and validation subsets we achieve similar performance to the previously trained weights file in completeness and purity calculated on the validation dataset but do so with less than one tenth of the amount of training data. The new weights also more consistently classify objects in the same way when used on unlabeled data compared to the old weights file. In the process of improving QuasarNET’s classification accuracy we discovered a systemic error in QuasarNET’s redshift estimation and used our findings to improve our understanding of QuasarNET’s redshifts.
@@ 1 Introduction
The Dark Energy Spectroscopic Instrument (DESI) is an ongoing spectroscopic survey that aims to observe over 40 million objects over the course of five years, with the goal of mapping the expansion history of the universe [1, 2]. DESI is mounted on the Mayall 4-meter telescope at Kitt Peak National Observatory and will map approximately a third of the night sky over the course of the survey [3, 4]. DESI takes an exposure of up to 5000 spectra at a time, a significant increase in data collection over previous spectroscopic surveys [5, 6, 7]. To date [8] approximately one million spectra were released publicly in the DESI Early Data Release (EDR) [9, 10], followed by more than 18 million spectra in Data Release 1 (DR1).
@@
QuasarNET originally consisted of four convolution layers followed by a fully connected layer. This fully connected layer was then connected to a series of fully connected layers of 26 outputs each. The number of fully connected layers is determined by how many quasar emission lines QuasarNET is trained on. Each output fully connected layer represents one line-finder n=unit. The first 13 outputs of each line-finder unit is a confidence value that the given emission line is present within that portion or “box” subset of the wavelength range, a coarse-grained estimate of the line position. The second 13 outputs are fine-grained estimates of line positions within each of the 13 boxes and are floating point values between -0.1 and 1.1, where 0 corresponds with the lower edge of the box and 1 corresponds with the upper edge. Classification of spectra using these outputs is done by selecting a confidence threshold and a number threshold and assigning a quasar classification to any spectra where more lines than the number threshold have confidence values above the confidence threshold. For example, in DESI any spectrum that has at least one emission line (number threshold of one) of confidence greater than 0.95 (confidence threshold of 0.95) in QuasarNET’s output will be classified as a quasar. Classification therefore only only uses the first 13 outputs of each line-finder unit, namely those that are associated with line confidence. QuasarNET also provides a redshift estimate for each spectrum, producing an independent redshift estimate for each emission line detected in the spectrum. The final reported redshift by QuasarNET is then the redshift of QuasarNET’s most confident line. Redshift estimation is done by using the combined estimated line position (coarse + fine) to estimate a central wavelength of an emission line, which is then converted to a redshift using the rest-frame wavelength associated with that line-finder unit. [21] provides a more in-depth explanation of how QuasarNET works.
@@
After outlining our modifications to QuasarNET, we outline the active learning methodology and its application to improving QuasarNET’s classification accuracy and purity when coupled with the aforementioned network changes. Ideally we would like a comparable amount of DESI data to train a new weights file to improve QuasarNET’s handling of DESI data. The sheer quantity of spectra observed by DESI, however, makes visual inspection of DESI data on a similar scale unfeasible, so by construction any training dataset comprised of DESI data will have significantly fewer spectra than that comprised of eBOSS data. To combat this reduced amount of training data we apply the active learning methodology to select an optimal set of data to be labeled that will provide the largest improvement in the network when trained. We couple this with a novel outlier rejection step designed to further ensure that the spectra we select to label provide adequate coverage of the possible data space. This methodology shows significant self improvement, with increase in both completeness and purity after each active learning step. The final produced DESI trained weights file shows similar performance to the previous eBOSS trained weights file, but uses a much smaller training dataset of approximately 6000 spectra, only 10% of the previous training sample.
On labeled DESI data our new weights file performs similarly to the previously trained weights file, with a marginal improvement in completeness and similar purity. We demonstrate this by splitting our selected sample into training and hold out validation sets. We then demonstrate the new QuasarNET’s performance on unlabeled data by comparing the classification and redshift estimates of repeated exposures of the same objects. In this test we further demonstrate additional benefits of our new dataset with improved repeat classification accuracy, with QuasarNET more consistently classifying objects the same way in multiple exposures.
@@ 3 Active Learning and Dataset Selection
During the first few months of the DESI survey a selection of deep tiles were visually inspected to quantify the performance of the automatic DESI classification pipeline [27, 28]. These efforts visually inspected approximately 16000 galaxy targets of various subtypes and approximately 8500 quasar targets. We use a subset of these spectra as an initial training dataset, keeping all the quasar specific tiles and a small subset of galaxies from the galaxy tiles. The small size of this dataset limited the performance of the resultant QuasarNET weights file.
To improve performance we sought to label additional DESI spectra to supplement the initial training dataset. Since the amount of human effort to label and classify spectra is non-trivial, we want to minimize the amount of new spectra labeled while maximizing the impact of those spectra on improving QuasarNET’s performance. We use the first few months of DESI main survey observations, over 3.6 million spectra, processed in the “Guadalupe” data reduction and released as a supplement to DESI’s DR1, as a bank of unlabeled spectra from which we will select spectra to label. In order to select which of these spectra to label and include in training we use the active learning algorithm. For an in depth and comprehensive exploration of active learning, its implementation, and its mathematics, we recommend [29].
Briefly, active learning uses an algorithm’s classification uncertainty to quantify which unlabeled data is most confusing to the network. This confusing data is then passed to a human expert who labels them for inclusion in a new training run. By continuously training on confusing data we can maximize the performance of the algorithm with the smallest amount of data, as we are continuously filling the network’s gaps in knowledge. When used with a classification algorithm active learning will tend to pick data points that lie along decision boundaries, i.e. those that have the highest uncertainty and thus the largest possibility of classification confusion.
@@
Once the SOM has removed any potential outlier spectra we select the 1000 objects with the highest entropy values to be visually inspected and reincorporated into the training dataset. We select 1000 spectra to balance the impact the new data has on the network against the speed at which we can visually inspect spectra. The active learning algorithm we perform can be summarized by the following steps:
  1. 1.
    Begin with a set of training data ($T$) and a set of unlabeled potential training data ($U$).
  2. 2.
    Bootstrap 200 unique subsets of $T$ to train 200 unique versions of QuasarNET.
  3. 3.
    Run all of $U$ through the bootstrapped QuasarNET versions, generating 200 unique classifications and redshift estimates of all spectra in $U$.
  4. 4.
    Use a SOM to remove any spectra from $U$ that are outliers, specifically any spectra in bins whose counts are below the 15th percentile of all bins.
  5. 5.
    De-duplicate spectra of the same objects in $U$, so that the potential dataset consists only of unique objects.
  6. 6.
    Select 1000 objects from the remainder of $U$ to be visually inspected and labeled.
  7. 7.
    Label and classify the 1000 selected objects.
  8. 8.
    Incorporate the newly labeled 1000 objects into $T$ and remove them from $U$.
After conclusion of step 8 the algorithm can be repeated as many times as desired. We refer to a single iteration of these steps as a “run”. In this work we perform two runs of active learning.
The output of a single run of the entire active learning algorithm is represented in Figure 3. We start with a potential dataset of 3.6 million spectra and apply the SOM based outlier rejection step to eliminate 15.5 thousand spectra from the pool of potential training data. We then apply a cut, removing repeat exposures of the same object to ensure that we have unique objects to visually inspect rather than multiple exposures of the same objects. This reduces the pool of spectra to just below 3 million unique objects.
In Figure 3 we plot a histogram of the entropy values, defined in Eq. 3.2, of the remaining spectra. A red line indicates the cutoff we selected, above which there are 1000 spectra that we request to be visually inspected and labeled. The plot shows two notable features: first, there is a sharp peak at zero, representing a large set of spectra for which there is high degree of confidence and agreement in the ensemble classification. Second, there is a sharp dropoff near 1 corresponding to the point at which a spectrum lies at the 50% confusion boundary between two classes, with no classifications in the third. Above this point in the histogram are spectra that show significant confusion between all three classes, rather than being either confused between only two classes or highly confident in one and weakly confident in the others.
@@ 4 Results
We ran the active learning pipeline twice, selecting 1000 spectra at each step to visually inspect and label. At the initial step we use 3761 DESI spectra from quasar dense tiles (“Deep-field VI” from Table 1 of [27]) plus an additional 839 DESI spectra from non-quasar dense tiles (a subset of all spectra in Table 1 of [28]) which we include to ensure that the dataset is not overly quasar dense. These 4600 DESI spectra were visually inspected and the results validated in [27, 28]. In the following section this dataset of 4600 spectra is named DESI VI. We then augment this dataset with an additional 15000 spectra from the previous eBOSS training dataset. We are motivated to supplement the DESI-only dataset for the first pass because the DESI only dataset does not fully cover the possible quasar space. We select only 15000 spectra to avoid overwhelming the DESI data when training the network; the additional spectra are designed to augment the quasar space covered by the DESI spectra without superseding any properties of the DESI spectra not present in the eBOSS data. 15000 eBOSS spectra ensures that roughly a quarter of the total dataset for this training run are DESI spectra. For our initial active learning run, $T$ thus has a size of $19600$ spectra.
For the second active learning step we use only the originally labeled DESI spectra plus the additional 1000 labeled spectra from the first active learning step to train the networks, excluding the supplemental eBOSS data. At this point we now consider it representative of the DESI quasar observation space, and believe that the active learning algorithm will select spectra to fill any deficiencies. This dataset of 5600 spectra (the initial 4600 + 1000 newly labeled spectra) will be referred to as DESI AL 1. In the second active learning run we will label an additional 1000 spectra. In the second visual inspection campaign we additionally investigated 100 high-entropy spectra that were rejected as outliers to verify that the SOM has identified outlier-like spectra. The combination of the initial 4600 spectra, the 2000 active learning spectra, and the 100 outlier spectra (for a total of 6700 spectra) will be referred to as DESI AL 2. We will also refer to a fourth dataset in this section, eBOSS DR12, comprised of the set of approximately 60,000 (eBOSS) spectra used to train the previous version of QuasarNET whose performance we are comparing to.
@@
For the visual inspection steps in the active learning pipeline we require that three separate experts inspect each spectra, assigning an estimated redshift and spectral classification as well as a “quality” rating that quantifies how confident the inspector is in their labels. Quality ratings are assigned from zero to four, with four the most confident, and while each expert makes their own determination of their own confidence they are broadly instructed that quality zero is reserved for if there is no data. Qualities one and two are generally for uncertain or insecure redshifts, and three and four are used for secure redshifts. Each spectra is then inspected by a “merger” who resolves discrepancies in the three expert’s labels and assigns a final merged quality rating. A more detailed description of this procedure is available in [27] and [28]. QuasarNET only uses spectra whose final merged quality is above 2.5.
@@
We present the results of this two-step procedure in three subsections. The first subsection includes a short discussion on the inspection of the rejected 100 spectra we visually inspected during the second active learning run, justifying their inclusion in the DESI AL 2 dataset. The second and third subsections outline our results: in the second we report results using the full truth dataset which we split into training and validation subsets. In the third and final subsection we test the final DESI-specific weights file, trained on the entirety of the truth table rather than a subset. The weights testing used the internal data release “Guadalupe”, which is included in DESI’s Data Release 1 (DR1), where we have no truth data but for which we do have a set of metrics that quantify this weights file’s improvement over the previous version. These metrics include consistency in classification and redshift uncertainty. In the third section we will also outline the new understanding of QuasarNET’s redshift estimation problems, which became apparent when using this newly trained weights file.
4.1 Investigation of High Confusion SOM-Rejected Outliers
Figure 4: An example of four representative high confusion spectra rejected by the SOM as outliers. Each plot shows the raw spectrum from the “Guadalupe” reduction as the lighter curve. The darker overplotted line in each plot is the spectrum smoothed with a Savgol filter with window size 75 pixels and polynomial order 3. The top (blue) spectrum is a high SNR and low redshift ($z\lesssim 0.4$ galaxy). The next (magenta) spectrum is a quasar with weak or no CIII emission which is expected around 6000 Å. The next (red) spectrum is a quasar with narrow CIV emission around 4800 Å. The bottom (orange) spectrum is an unusually high SNR quasar. The top three spectra all show an unusual high flux pixel near 5600 Ådue to the sky subtraction and calibration.
In this section we will briefly review the visual inspection of the 100 outlier spectra included in the second visual inspection campaign. These 100 spectra were rejected by the SOM and were investigated to determine if they were correctly rejected as outliers. These 100 spectra were the highest confusion spectra in the second active learning run that were rejected by the SOM, and were visually inspected using the same procedure as the rest of the (non outlier) data.
The SOM rejects spectra if there’s too few other similar spectra in the dataset, so without visual inspection we cannot determine exactly why a spectrum was rejected. Most if not all of the 100 spectra visually inspected are correctly identified as outliers. Broadly many of these outliers are outliers fall into two categories: data quality outliers and object type outliers. A few representative spectra are shown in Figure 4. These four representative spectra exhibit characteristics present in all of the 100 inspected outliers. The top three panels (blue, purple and red) show object type outliers: a low redshift galaxy with strong emission lines, a quasar with weak to no CIII emission, and a quasar with narrow CIV absorption respectively. The last spectrum is of note because no broad CIV absorption was found in the sample of rejected spectra, only those with narrow absorption. The bottom (orange) panel shows a data quality outlier: the quasar has an unusually high SNR and strong SIV. Even among outliers this is comparatively rare, most of the data quality outliers are spectra that have such low SNR that identifying spectral features is impossible, if they exist at all.
From this sample we concluded that the SOM is accurately rejecting outliers, even if some of the outliers may be surprising. All four of these representative spectra also represent areas of high confusion for QuasarNET: the first three are unusual objects with rare or unusual spectral features, the last is so high SNR that QuasarNET disagrees on line assignment (leading to redshift confusion, this spectrum lies along the HIGHZ/LOWZ confusion boundary). Since we were able to generate secure, high quality visually inspected redshifts for these 100 spectra, we opted to fold them into the DESI AL 2 dataset.
@@ 4.2 Results on Labeled Data
In this section we train a weights file using a subset of the full truth table (DESI AL 2), which includes the original set of visually inspected quasar tiles (3761 spectra), the additional subset of other galaxy dense visually inspected tiles (839 spectra), and the results of two active learning runs (2100 spectra). The total number of spectra in the truth table is 6700, although after training we discovered 2 spectra whose data processing was corrupted that were affecting training and removed them for a final total of 6698 labeled spectra. These two spectra were both visually inspected with high confidence based on the uncorrupted data, but the presence of data processing corruption on approximately 1/3rd of the wavelength grid in these two spectra made them unsuitable for use in QuasarNET. We split this dataset into two subsets such that 70% of the data is used for training and the other 30% is used for testing and validation. The validation dataset is comprised of 2010 spectra, the remaining 6790 are available for use as training data.
In order to validate that the inclusion of the active learning spectra has improved the network’s performance we compare this training to three other training datasets: DESI VI, DESI AL 1, and eBOSS DR12. The results of these four runs are shown in Figure 5. DESI VI is shown in blue, DESI AL 1 is in magenta, DESI AL 2 is in red and eBOSS DR12 is in orange. For each of the other training schemes we use the same validation dataset (the subset of DESI AL 2 reserved for validation) excluding from the given training dataset any spectra that appear in the validation data. The numbers in the plot represent the number of spectra used to train that network after removing spectra in the validation dataset and after making the QuasarNET visual inspection quality cuts. For this plot, we use a confidence threshold of 0.95, declaring that QuasarNET identifies a quasar if it has at least one line with a confidence of more than 0.95.
@@
[…]
As validated in the plot, DESI AL 2 has a significant improvement in completeness at approximately the same (although marginally better) purity at each epoch compared to DESI AL 1. DESI AL 1 has a more impressive gain in completeness over the original DESI VI, although it does suffer from decreased purity. Compared to eBOSS DR12, our final run, DESI AL 2, is comparable in completeness and purity, but uses less than a tenth as much data.
The training curves shown in Figure 5 show some oscillation and are not smooth. This is because the training loss function used to train QuasarNET tries to minimize the distance between the predicted line position and the true line position, equally weighting the coarse and fine grained estimates (for more details see Equation 1 of [21]). Classification uses only the coarse grained confidence in line position, while redshift estimates use the both the coarse and fine grained output. At some epochs the completeness or purity may fluctuate as confidence values change, but at the same epoch the redshift estimation may improve in accuracy. We only investigate the purity and completeness here as we are only concerned with classification accuracy as opposed to redshift accuracy.
@@
It is possible that a larger dataset could explain some or all of the improvement at each active learning step. In order to validate that our improvement is more significant than dataset size alone, we bootstrap each of the two earlier runs, DESI VI and DESI AL 1, to have the same amount of data as the final assembled dataset. We perform 10 unique bootstraps of each dataset, for 20 total unique datasets of the same size as the final dataset, and then train for 100 epochs with the same parameters as the previous test. This bootstrapping simulates naively adding additional spectra to each dataset drawn from the same distribution as that dataset.
The results of this test are shown in Figure 6. For clarity we plot only the completeness, although purity follows a similar trend. The solid lines represent the same completeness curves as plotted in Figure 5. In dashed blue and dashed magenta are the 10 bootstrap runs for the DESI VI and DESI AL 1 datasets respectively. Although the increased amount of data does have some effect on the training it does not significantly alter the final best completeness values. This is not unexpected but it does validate that the active learning methodology has specifically selected spectra that expand the quasar space covered by the training dataset, rather than simply sampling similar quasars to those already covered by the training data.
@@
Figure 7: Plot of completeness versus purity for a variety of confidence thresholds ranging from 0 to 1 using the final produced weights file for each of the four datasets. At each confidence threshold we compute quasar classifications and the corresponding completeness and purity on the validation dataset given those classifications. Colors match those of Fig. 5. A large dot on each curve (below the bottom of the plot for the DESI VI dataset) denotes the values at the standard 0.95 confidence threshold. At a confidence threshold of 0.95, DESI VI 2 outperforms the eBOSS weights file in completeness with a similar purity, although there are some threshold values at which the opposite is true.
For our active learning pipeline we used a confidence threshold of 0.95, meaning that any spectrum with at least one line with a confidence more than 0.95 was considered a quasar. This motivated the choice used in the loss curves in Figure 5 and Figure 6, where the purity and completeness per epoch were calculated using this same confidence threshold.
It is also informative to explore how the choice of confidence threshold changes the final results of the network. In Figure 7 we plot the purity and completeness against each other for the final networks trained on each of the four datasets. The curves are computed by varying the confidence threshold used for classification between 0 and 1, showing the trade off between improved completeness and lower purity at lower thresholds and the opposite at higher thresholds. For example, with a very low confidence threshold ($\sim 0.1$), the sample will be very complete but relatively impure as almost everything will be classified a quasar. A high threshold ($\sim 0.99$) will generate a very pure but incomplete sample as only high confidence spectra are labeled quasars. Increasing the confidence threshold has the effect of moving along a curve from the top left of the figure (low threshold) to the bottom right (high threshold). A perfect classifier is at the top right corner of the figure (100% pure and 100% complete).
@@
Large dots in the figure indicate the purity and completeness of each network at DESI’s nominal confidence threshold of 0.95. At the nominal threshold of 0.95 we have succeeded in improving the completeness of the final weights file, with almost identical purity. Notably this effect does not hold true at all confidence thresholds. At lower thresholds the eBOSS DR12 weights file exceeds our newly trained weights files in completeness with similar or better purity. This is due to our use of this confidence threshold when determining classification for active learning, as we only pick spectra that are confused between classes when the constituent networks of the ensemble are very confident in their own classifications. It is likely that if we had used a different threshold we would see our network outperforming the eBOSS one at that specific threshold, as we do with a threshold of 0.95.